# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v7

@@ 7 Interior Degrees Doubling
Lemma 7.1.
Indistinguishable Regular Interior Nodes Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER. We are attempting to prevent a degree-doubling node. If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container. Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of containers. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures. This means any node will have its interior degree doubled under the automorphism $\phi$. ∎
We still need to prove that for each container there exists a node whose interior degree doubles. We have seen this in the example 5.5, but those could be too simple to observe any real patterns. In similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2. However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
@@
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER. We are attempting to prevent a degree-doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
@@
We prove this by strong induction on the size of the graph. Base Case: In Lemmas 4.1 (Minimum Out-Degree $<$ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
@@
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.), there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
@@
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples. The questions of whether this was always true or whether back arcs like lemma 6.3 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container we can simply run the conjecture on that container. The difference is that with lemma 7.2 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container. Then with lemma 7.1 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have their interior degree double.
@@ 8 Container Density
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP). Then the nodes in $C_{i}$ will result in regular exterior degrees.
@@
Because Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs) both hold, we can apply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
This decreasing of the exteriors (and thus the containers) is stopped at the container $C_{i}$. This is where our assumption of the different exterior sizes comes into play. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
@@
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 5.6 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
@@
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors, and thus $C_{i}$ must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$.
@@
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$. Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1 (Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
@@
In Section 5 (Theoretical Foundation), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the chosen minimum-degree node. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from the minimum degree node entirely. Still, the use of induction provides an important method, particularly for graphs where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
@@
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a parent of $x$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$.
@@
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors. Once again, the minimum degree node is the driving force, the catalyst, behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
Lemma 8.4.
Regular Graph All Nodes Have Equal Out-Degree Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
Proof.
For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs. We have already shown that back arcs cannot exist. If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$. While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
This more rapid reduction in the interior degree leads to container density faster, which would simply lead to a degree doubling node faster. To avoid this, overall regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph. ∎
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the decreasing sequence property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.4 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
@@ 9 Transitive Triangles and Seymour Diamonds
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm. Transitive triangles represent a common situation with our containers. A parent node $x$ has two children $y$ and $u$ that are also connected by an arc. The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double. We have already shown through 8.3 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
@@
Seymour Diamonds from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
@@
Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
@@
None of these cases are a deterrent to the execution of the DSP algorithm.
Lemma 9.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds. These diamonds are constructed to be presented in our containers. As a result, we simply select an interior node within the container and continue. This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example 4.3. This is also what we saw in our social media example 5.5 in section 5. This is also what we saw in our regular graph example in 8.1 in section 8.
@@
Interaction of Minimum Degree Nodes and Transitive Triangles
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $\delta$. Consider a transitive triangle formed by nodes $u$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(u)=0$.
Then the following cases are possible:
  • $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$:
@@
  • Case 1: $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$: Both nodes $y$ and $u$ out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has $\delta$ out-neighbors that are not in the triangle. Since we assumed that $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$, we know that $y$ cannot connect to all not $u$’s neighbors since $y$ has a neighbor in the triangle ($u$). This node that is $u$’s neighbor, but not $y$’s neighbor, is a candidate to extend the path.
  • Case 2: $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree $\delta$. The node $u$ has out-degree greater than $\delta$. Again, the triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has more than $\delta$ out-neighbors that are not in the triangle. We assumed that $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$. We know that $y$ cannot connect to all not $u$’s neighbors since $d^{+}(u)>d^{+}(y)$. We can then select one of these nodes that is a neighbor of $u$, but not a neighbor of $y$ to be a candidate to extend the path.
  • Case 3: If $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. The node $y$ has more than $\delta$ out-neighbors that are not in the triangle. The node $u$ has exactly $\delta$ out-neighbors that are not in the triangle. The maximum number of neighbors that $y$ and $u$ can share is $\delta$, that being $u$’s out-degree. Then we can select one of the nodes that is a neighbor of $y$, but not a neighbor of $u$ to be a candidate to extend the path.
  • Case 4: If $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree greater than $\delta$. We have a few cases because we have to remember that in this data structure the purpose of our containers and the interior nodes was to prevent parent nodes from having their degrees double. The case where both nodes in a transitive triangle have more than $\delta$ out-neighbors means that $x$ has up to $2*\delta$ out-neighbors, which would cause its degree to double. Thus, to prevent $x$’s degree from doubling, there must be some overlap between the out-neighbors of $y$ and the out-neighbors of $u$. These two sets cannot be equal to one another, though. The most overlap we can have is $N^{+}(u)\subsetneq N^{+}(y)$ or $N^{+}(y)\subsetneq N^{+}(u)$. Both of these situations leave room for a node in the larger set that is not in the smaller set as a candidate to extend the path. As we allow for less overlap, the number of candidate nodes to extend the path grows. We can see that there will always be a node in the larger set to extend the path.

### v8

@@ 8 Interior Degrees Doubling
Example 8.1.
[…]
Figure 9: Illustration of the proof of 8.1 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
We still need to prove that for each container there exists a node whose interior degree doubles. We have seen this in the example 6.3, but those could be too simple to observe any real patterns. In a similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2. However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
Example 8.2.
@@
Interior Degrees Double Let $G$ be an oriented graph. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
@@
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree $<$ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
@@
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph.
By the Induction Hypothesis (I.H.), there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$.
We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$.
We can define $H_{3}=ext(y_{1},y_{2})$.
Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
@@
Lemma 8.1 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. We were able to set up a GLOVER data structure inside this local version starting with a minimum degree node of the container. We can then form local containers that are each their own respective distance from the minimum degree node. By strong induction, we have that each of these containers will have a node whose interior degree doubles. Then by the weak decreasing exteriors lemma, we have a decreasing set of exteriors inside this container. When this sequence gets down to size two, that node will have all exterior neighbors. Therefore, a parent of that node will have its degree doubled.
By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples. The question is whether this was always true or whether back arcs like lemma 7.3 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container, we can simply run the conjecture on that container. The difference is that with lemma 8.1 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container. Then with lemma 9.2 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have its interior degree doubled.
Example 8.3.
[…]
Figure 10: Illustration of the proof of 8.1 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
@@ 9 Container Density
Regular Exterior Degrees Let $G$ be an oriented graph. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP). Then the nodes in $C_{i}$ will result in regular exterior degrees.
@@
Because Lemmas 8.1 (Interior Degrees Doubling) and 7.5 (No Back Arcs) both hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
This decreasing of the exteriors (and thus the containers) is stopped at the container $C_{i}$. This is where our assumption of the different exterior sizes comes into play. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node parent node in $C-{i-2}$ can have its exterior degree double. By 8.1, we know that this node has its interior degree double as well.
@@
This proof begins with two nodes in the same container. By that assumption and the regularity lemma on interior degrees, we know that these nodes have the same interior degree. Our next assumption is that these two nodes have different exterior degrees. We simply want to observe what effect that has on a parent of these nodes. The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by the weak Decreasing Exteriors Lemma. The container $C_{i+1}$ is now the same size as $C_{i}$. This means that a node in $C_{i-1}$ will have their exterior degree double.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
@@
Indistinguishable Regular Interior Nodes Suppose we have a regular container $C_{i}$ of size $n$ of an oriented graph $G$. If a node $v_{i}$’s interior degree doubles, then every other node $v_{j}$ in the same container as $v_{i}$ will also have its interior degree double.
@@
Let $C_{i}$ be a regular container, and Let $v_{i}$ be a node in this container. Further, let $v_{p}$ be a parent of $v_{i}$. Then $\text{int}(v_{p},v_{i})$ doubles in $C_{i}$ by 8.1 (Interior Degrees Doubling). In a regular container, all nodes have the same interior by 7.1 (Regular Interior Degrees) and exterior degrees by 9.1 (Regular Exterior Degrees). They also have the same parents because of containers, also by 9.1. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve this degree structure. This means any node will have its interior and exterior degree doubled under the automorphism $\phi$. Therefore, the node $v_{j}$ which is mapped to by $\phi$ will have its interior degree doubled. ∎
Now that we have a collection of nodes inside a container that all have equal interior and exterior degrees, as well as the same parents. There becomes a question of how we can distinguish between these nodes. The truth is that we cannot distinguish between them. As a result, we are led to conclude that if one node in a container has its interior degree double, then all nodes in that container will also have this property.
Lemma 9.3.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph, where $v_{0}$ is a minimum degree node in $G$. Then, for any container $C_{i}$, the size of the exterior set leading to container $C_{i+1}$ is bounded by
|$|\{y\in\text{ext}(C_{i},z)\text{ s.t. }y\in C_{i+1}|\leq d^{+}(v_{0})-i$
Proof.
We will prove the statement by induction on the container index $i$.
Base Case ($i=1$):
Container $C_{1}$ consists of the out-neighbors of $v_{0}$, which are distance 1. By definition, the exterior set at distance 2 is composed of the neighbors of nodes in $C_{1}$. Since $v_{0}$ has a degree of at least 1, we obtain:
$|{y\in\text{ext}(C_{1},z)\text{ s.t. }y\in C_{i}}\leq d^{+}(v_{0})-1.$
Inductive Hypothesis:
Assume that for some $i$, the bound holds for container $C_{i}$, i.e.,.,
|$|\{y\in\text{ext}(C_{i},z)\}\text{ s.t. }y\in C_{i+1}\}\leq d^{+}(v_{0})-i.$
Inductive Step:
Consider container $C_{i+1}$. These are the out-neighbors of nodes in $C_{i}$. Each node in $C_{i}$ satisfies the Decreasing Sequence Property (DSP), its second neighborhood is smaller than its first neighborhood. This implies that the number of new nodes in $C_{i+1}$ must be strictly smaller than the size of $C_{i}$.
Alternatively, we could apply the 6.4 Weak Decreasing Exteriors Lemma To $C_{i}$. We observe that since there are no back arcs and by 8.1 (Interior Degrees Doubling) there exists a node whose interior degree doubles, it follows that $C_{i+1}$ is strictly smaller than $C_{i}$. Thus, the size of the exterior set satisfies:
|$|\{y\in ext(C_{i+1},z)\text{ s.t. }y\in(C_{i+2}\}|\leq d^{+}(v_{0})-(i+1).$
By induction, the bound holds for all container levels $i$
@@
In Section 5 (Building a Data Structure), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the chosen minimum-degree node. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from $v_{0}$ entirely. Still, the use of induction provides an important method, particularly for graphs, where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose degree doubles. The difference is that that assumption does not need to be explicitly stated because we have already proved that there is a node whose interior degree doubles in each container. Similarly, this lemma requires there to be no back arcs, like the weak decreasing exteriors lemma. We’ve already proved that back arcs cannot exist without creating a degree doubling node, though. As a result, we do not need to state that assumption.
The proof then simply iteratively executes lemma 6.4 (Weak Decreasing Exteriors Lemma), counting the size of the exteriors as we extend our distance from $v_{0}$. This can be summed up by the formula $|C_{i+i}|<d^{+}(v_{0})-i$.
@@
Interior Degrees Lower Bound Let $G$ be an oriented graph. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a parent of $x$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$.
@@
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors. Once again, the minimum degree node $v_{0}$ is the driving force, the catalyst, behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
Once again, we see a proof with not much that needs to be assumed. The only real assumption here is that we have a container $C_{i}$ in the oriented graph $G$ that is distance $i$ from $v_{0}$. Lemma 9.3 (Strong Decreasing Exteriors Lemma) tells us that $|C_{i}|<d^{+}(v_{0})-i$. By the regularity of exterior degrees, we know they all nodes in $C_{i}$ map to $C_{i+1}$. With the absence of back arcs, the only way for nodes in $C_{i}$ to meet the minimum degree requirement is to add more interior ones. Each node would need at least $i$ interior arcs to satisfy this inequality.
@@ 10 Transitive Triangles and Seymour Diamonds
Lemma 10.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm. Transitive triangles represent a common situation with our containers. A parent node $x$ has two children $y$ and $u$ that are also connected by an arc. The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double. We have already shown through 9.4 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
@@
Seymour Diamonds from Containers Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
@@
Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph, where all nodes have the Decreasing Sequence Property (DSP). We are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
@@
None of these cases are a deterrent to the execution of the DSP algorithm. ∎
Lemma 10.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds. These diamonds are constructed to be presented in our containers. As a result, we simply select an interior node within the container and continue. This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example, 4.3. This is also what we saw in our social media example 6.3 in section 5. This is also what we saw in our regular graph example in 9.1 in section 8.
@@
Node in GLOVER’s Transitive Triangles Suppose $G$ is an oriented graph. Let $v_{0}$ be a minimum degree node and let $u$ be a node in container $C_{i}$. Then $u$ is in three possible locations of two different transitive triangles, which do not hurt the algorithm.
@@
Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
Scenario 1 is that $u$ is a parent to children $v$ and $w$ who are both in $C_{i+1}$, where $v,w)\in G$. Then $(u,v)$ and $(u,w)$ are exterior arcs, and $(v,w)$ is an interior arc.
Scenario 2 is that $u,v$ are children of $x$ and $u$ is a child of $v$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(v,u)$ is an interior arc.
Scenario 3 is that $u,v$ are children of $x$ and $v$ is a child of $u$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(u,v)$ is an interior arc.
In every scenario, we have shown in preceding sections how to proceed through these situations. So these do not hinder our algorithm. ∎